What is the molarity of a solution made by diluting (2.1067x10^1) mL of 18.7 M ammonia with enough water to form (2.07x10^2) mL of solution?
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 18.7 M

V1 = 21.067 mL

V2 = 207 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (18.7*21.067)/207

M2 = 1.9032 M

Answer: 1.90 M